OpenStudy (anonymous):

The following are the schedule of matches in a chess tournament : Day1: Jayson vs. Jethro Day2: Charlene vs. Jethro Day 3 : Charlene vs Jayson What is the probability that Charlene wins both her games?

OpenStudy (anonymous):

i cant understand the question...

OpenStudy (anonymous):

Asking for probablity, based on the sched.?

Parth (parthkohli):

There are 2 possibilities. She will win or she won't win, so we'll multiply the probabilities. \(\Large \color{MidnightBlue}{\Rightarrow {1 \over 2} \times {1 \over 2} }\)

Parth (parthkohli):

Just asking, but is there a possibility of a draw?

OpenStudy (anonymous):

the question isnt complete i guess....

Parth (parthkohli):

If there's a possibility for a draw then it'd be 1/3 * 1/3 = 1/9

OpenStudy (anonymous):

The following are the schedule of matches in a chess tournament : Day1: Jayson vs. Jethro Day2: Charlene vs. Jethro Day 3 : Charlene vs Jayson What is the probability that Charlene wins both her games? each player has equal chances of winning as losing.

Parth (parthkohli):

So then it'd be 1/2 * 1/2

OpenStudy (anonymous):

its 1/4 then ya?

Parth (parthkohli):

1/4 probability

Parth (parthkohli):

Yes

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

owww. got it :)

OpenStudy (anonymous):

if each player has equal chances of winning as losing, \[0.5\times0.5\]=1/4

OpenStudy (anonymous):

Thanks! :)))) HOw about this? What is the probability that Jayson wins one and loses one?

Parth (parthkohli):

Exactly P(winning) = 1/2 2 games, so 1/2 * 1/2 = 1/4

Parth (parthkohli):

1/2 of winning 1/2 of losing 1/2 * 1/2 = 1/4

OpenStudy (anonymous):

\[2\times0.5\times0.5=1/2\]

Parth (parthkohli):

Why the 2?

Parth (parthkohli):

WW WL LW LL 2/4 1/2 I see

Parth (parthkohli):

Correct, 1/2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

What is the probability that Jethro wins both his games and Charlene wins both her games?

Parth (parthkohli):

1/4 is the probability for both of winning. 1/4 * 1/4 1/16

OpenStudy (anonymous):

\[1/4\times1/4=1/16\]

Parth (parthkohli):

No no no.....

OpenStudy (anonymous):

why?

Parth (parthkohli):

It's 0 because one has to lose in the match they'll play.

Parth (parthkohli):

When they match, one of them HAS to lose.

OpenStudy (anonymous):

yes.0

OpenStudy (anonymous):

Genius, both of you! :) Thanks!!!!! :D

Parth (parthkohli):

Hehe lol

OpenStudy (anonymous):

hehe,i am wrong

OpenStudy (anonymous):

still, you stand corrected ;)

OpenStudy (anonymous):

Lol, 5 gum eh? XD

OpenStudy (anonymous):

...gum...

OpenStudy (anonymous):

Hahaha. Just ride XD

OpenStudy (anonymous):

xd....ok...